Which of the following statements are true and
which are false ? Give reasons for your
answers.
(a) Any square matrix with real entries is either symmetric or skew-symmetric or a linear combination of such matrices.
(b) If a linear operator has an eigenvalue 0,
then it cannot be one-one.
(c) If f : V --> K is a non-zero linear functional
and V a vector space of dimension n, then
there are n - 1 linearly independent vectors
v belongs to V such that f(v) = 0.
(d) Every binary operation on Rn is
commutative, for all n belongs to N.
(e) IAdj (A)I = IAI for all A belongs Mn(R).
1
Expert's answer
2020-11-25T16:33:45-0500

(a) Any square matrix with real entries is either symmetric or skew-symmetric or a linear combination of such matrices.

True


Let "A+A'" be a symmetric matrix "A-A'" be a skew matrix

Let "B=A+A'" then "B'=(A+A')'"

"\implies A'+(A')'\ (as\ (A+B)'=A'+B')\\
\implies A'+A (as\ (A')'=A)\\
\implies A+A' (as\ (A+B= B+A)+B\\"

Therefore "B=A+A'" is a symmetric matrix


Now let "C=A-A'"

"C'=(A-A)'=A-(A)'\ (Why?)\\
C'=A'-A\ (Why?)\\
\implies -(A-A')=-C"


Hence, "A-A'" is a skew-symmetric matrix

If "A" is a square matrix in "Mn(C)" then


"A=\frac12 (A+A')+\frac12(A-A')"

(b) If a linear operator has an eigenvalue 0, then it cannot be one-one.

False


Let "V" be an "n" dimensional vector space over an algebraically closed field F. Let x be a non-zero vector in V. If "Tx= 0" , then 0 is an eigenvalue. Assume that "Tx \ne 0" . Then we known that "\{x,Tx,...,Tnx\}" is linearly dependent, so that there exist "\u03b10,\u03b11,...,\u03b1k" in F with "k\{1,...,n\}" such that "ak \neq 0" and



"\u03b1_0x+\u03b1_1Tx+\u00b7\u00b7\u00b7+\u03b1_kT^kx= 0,"

i.e.,


"(\u03b1_0I+\u03b1_1T+\u00b7\u00b7\u00b7+\u03b1_kT^k)x= 0."

Thus,


"p(T)x= 0,"

where "p(t) :=\u03b1_0+\u03b1_1t+\u00b7\u00b7\u00b7+\u03b1_kt^k" . By fundamental theorem of algebra, there exist "\u03bb_1,...,\u03bb_k" in F such that


"p(t) =\u03b1_k(t\u2212\u03bb_1)\u00b7\u00b7\u00b7(t\u2212\u03bb_k)"

Since "p(T)x= 0" , we have


"\u03b1_k(T\u2212\u03bb_1I)\u00b7\u00b7\u00b7(T\u2212\u03bb_kI)x."

This shows that at least one of "T\u2212\u03bb_1I,...,T\u2212\u03bb_kI" is not one-one.

(c) If f : V --> K is a non-zero linear functional and V a vector space of dimension n, then there are n - 1 linearly independent vectors v belongs to V such that f(v) = 0.

False


By definition,

Suppose V is a vector space of dimension n, and that "v_1, v_2,...,v_n" are vectors of V. The set of vectors "\{v_1, v_2,...,v_n\}" is linearly independent.


If "f_1v_1+f_2v_2+...+f_nv_n=0" for some "f_1,f_2, ..., f_n \in F" where at least one of "f_1,f_2, ..., f_n" is a non-zero. This implies that there are n linearly independent vectors of V as opposed to n-1


(d) Every binary operation on Rn is commutative, for all n belongs to N.

True


Let∗ be a binary operation on a set Rn. Given any three elements "m,n" and "p" of a set Rn, the binary operation, applied to the elements m∗ n and p of Rn ,yields an element (m∗n)∗p of Rn, and, applied to the elements m and n∗p of Rn, yields an element m∗(n∗p) of Rn.


(e) IAdj (A)I = IAI for all A belongs Mn(R).

False


Every invertible matrix is a product of some elementary matrices. This

is seen by applications of a series of elementary row operations to the matrix

to get the identity matrix I. When matrix A is invertible, the inverse can

be found by the adjoint, the formula


"A^{-1}=\frac{1}{|A|}adj(A)"

It is easy to see that "|kA| = (k^n)|A|" where k is a constant and n is the order of the square matrix A.


We know that "|A^{-1}|\ is \frac{1}{|A|}" .Since "A^{-1}" is "\frac{1}{|A|} adj (A)", we get "|A^{-1}| = (\frac{1}{|A|})^n |adj(A)|" which implies "\frac{|A|^{(n)}}{|A|}" is "|adj(A)|" .


So "|A|^{(n-1)} = |adj(A)|" .